-3+-4=
-4x0=
0 divided by-7
4/5 divided by 1/5 =

Answer by Mathtut(3670)

-3+-4=-7
-4x0=0
0 divided by-7=0
4/5 divided by 1/5 =(4/5)*(5/1)=4